# Zero-Shot Machine-Generated Text Detection Using Mixture of Large Language Models

### v1

@@ 6 Summary and Discussion
As demonstrated, our RSA method is effectively harnessing the strength of the models in the ensemble, offering a scalable solution adaptable to various languages and domains, and also evolving over time. Yet, it displayed varying levels of performance across different text domains, suggesting a need for further research into the domain-adapted models on domain-specific ensemble strategies to improve performance in niche areas (e.g., the academic dataset we used). Furthermore, given that RSA-(max) sometimes outperforms RSA-(avg) and at other times does not, it is crucial to enhance the robustness of our approach. We must develop information-theoretic tools for understanding and selecting the most useful models while filtering out the less effective (or toxic) ones, as demonstrated in the case of the unigram model.
7 Limitations
Our method has proven to be quite effective in detecting content generated by a variety of diverse generators. We however acknowledge a series of shortcomings that will need to be addressed in future work. First, RSA is compute-intensive, as it requires, for each input token, to (a) compute the forward pass for $N$ LLMs, all having significant memory footprints and compute requirements; (b) to run the Blahut-Arimoto algorithm. As a result, each document in our implementation requires 10-20s per document, using a machine equipped with 4 GPUs (V100, 32G). A second limitation of our current approach is the requirement for each detector model to share the same event space, hence the same underlying tokenization. There is no simple solution available in the literature to this problem, which is also encountered in teacher-student knowledge distillation approaches (Raman et al., 2023). Finally, while we have tried to consider multiple generators and datasets for artificial text detection, our benchmarks are still heavily biased toward specific textual genres and languages, in particular clean and relatively long English texts, that correspond to “pure” generations. Testing our system also on other types of data, arguably more challenging, using new benchmarks such as M4GT-bench (Wang et al., 2024c), will help strengthen our robustness claims.
@@ Appendix A Proof of Proposition 1
where the minimum is taken over all the possible distributions $q\in\mathcal{P}(\Omega)$, representing the expected value of regret of $q$ w.r.t. the worst-case distribution over $\mu\in\mathcal{P}(\mathcal{M})$. Notice that this is equivalent to the average worst-case regret Barron et al. (1998); Silva and Piantanida (2022). The equality in (12) holds by noticing that
@@
is convex for every $(\mu,q)$, respectively. Then, by classical min-max theorem von Neumann (1928), we have that (15) holds.
Finally, it reminds to show that
@@
We begin by showing that
@@ B.1 Algorithm
Now if we denote the cardinality $|\mathbb{M}|=M$, $|\Omega|=N$, then $p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$ is an $M\times N$ matrix, which we denote the $i$-th row, $j$-th column entry by $w_{ij}$. For the case of channel capacity, the algorithm was introduced in Arimoto (1972); Blahut (1972) to solve (20). They both found the following expression for the capacity of a discrete channel with channel law $w_{ij}$:
@@
Then, using the theory of optimization, specifically coordinate descent, it has been showed that the sequence indeed converges to the required maximum. That is,
@@ B.2 Computational complexity
Appendix C Binocular Results
C.1 Binoculars score
@@
$B(\mathbf{w})\triangleq\frac{\displaystyle\displaystyle\sum_{t=1}^{T}\sum_{y_{t}\in\Omega}\mathbbm{1}[y_{t}=w_{t}]\mathcal{L}_{\theta}(y_{t}|m,\mathbf{y}_{<t})}{\displaystyle\sum_{t=1}^{T}\sum_{y_{t}\in\Omega}\displaystyle p_{\theta}(y_{t}|m^{\prime},\mathbf{w}_{<t})\mathcal{L}_{\theta}(y_{t}|m,\mathbf{y}_{<t})},$ (23)
@@
where the choices of the LLMs $p_{\theta}(\cdot|m,\mathbf{w}_{<t})\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ and $p_{\theta}(\cdot|m^{\prime},\mathbf{w}_{<t})$ $\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ are fundamentally important and have to be empirically selected. Indeed, this represents the main weakness of this score since in practice the best choice of the pair of LLMs $(m,m^{\prime})$ may not be distribution free.
C.2 Numerical results
We further develop on our Binoculars regeneration using Llama-2-7b on table 4, where it can be seen that the log-probabilities of distribution $q^{\star}(4)$ are as good a discriminator of generated texts as the average perplexity obtained with all 4 of our models. We believe this is because Llama-2-7b chat is almost always the best predictor for those samples. This behaviour changes on the Ghostbuster dataset when another instruct model, phi-3-mini-4k-instruct, is added to our ensemble (the reader is referred to 7 for further details). In Table 6, we also report the True Positive Rates at 5% False positive rate. We can see that the perplexities of our models are not good enough discriminators. RSA displays the most robustness across all three generators and domains. Note that in our case, a positive sample is a generated one, as the original Binoculars dataset has more generated samples than human ones, whereas our complementary generations keep the ratio 1:1.
Figure 1 shows the result of the follow-up study mentioned in Section 4.1. When computing the Blahut-Arimoto algorithm on the Pubmed dataset with our default RSA-4 configuration, the weight distribution is as shown on Figure 1(a). We applied a truncated scale to more effectively illustrate the discussed phenomena. This adjustment was necessary due to the high incidence of near-zero values, which would otherwise distort the visual representation of the data. When modifying the Llama-2-7b logits to match the sampling parameters used during generation, the algorithm gives much more importance to the aforementioned model, confirming that this modification suits the distribution of the data better.
Pubmed CNN CC_news
Original Regen-L Original Regen-L Original Regen-L
PPL oracle 0.823 0.690 0.598 1.000 0.617 0.632
PPL average 0.808 0.637 0.566 0.999 0.584 0.569
$q^{\star}$ (4) log-probs 0.807 0.641 0.566 1.000 0.579 0.568
Binocular score 0.988 0.633 0.995 1.000 0.979 0.578
RSA-4 (avg) 0.992 0.887 0.993 0.999 0.971 0.854
RSA-4 (max) 0.979 0.904 0.992 0.993 0.948 0.861
RSA-5 (+phi) 0.992 0.885 0.988 0.998 0.950 0.815
RSA-5 (+unigram) 0.986 0.846 0.960 1.000 0.898 0.766
Table 4: More results the Binoculars Dataset (AUC scores).
Pubmed CNN CC_news Average
Original Regen-L Regen-M Original Regen-L Regen-M Original Regen-L Regen-M
PPL TowerBase-7b 0.798 0.631 0.806 0.557 0.999 0.639 0.582 0.573 0.669 0.695
PPL TowerBase-13b 0.823 0.528 0.801 0.561 0.999 0.641 0.578 0.439 0.652 0.669
PPL Llama-2-7b-chat 0.817 0.690 0.843 0.598 1.000 0.716 0.617 0.632 0.718 0.737
PPL Llama-2-7b 0.786 0.676 0.820 0.544 1.000 0.704 0.556 0.617 0.693 0.711
$q^{\star}$ log-probs 0.807 0.641 0.836 0.566 1.000 0.727 0.579 0.568 0.703 0.714
Binocular score 0.988 0.633 0.892 0.995 1.000 0.956 0.979 0.578 0.922 0.883
RSA-4 (avg) 0.992 0.887 0.961 0.993 0.999 0.971 0.971 0.854 0.940 0.952
RSA-4 (max) 0.979 0.904 0.960 0.992 0.993 0.987 0.948 0.861 0.943 0.952
Table 5: Comparing results on Mistral-7B (Regen-M) and Llama-2-7B (Regen-L) generations (AUC scores).
Pubmed CNN CC_news Average
Original Regen-L Regen-M Original Regen-L Regen-M Original Regen-L Regen-M
@@ C.2 Numerical results
Binocular score 0.942 0.278 0.683 0.975 1.000 0.835 0.924 0.195 0.561 0.710
RSA-4 (avg) 0.963 0.615 0.814 0.971 1.000 0.898 0.868 0.552 0.730 0.823
RSA-4 (max) 0.904 0.660 0.800 0.964 0.967 0.954 0.782 0.538 0.742 0.812
Table 6: Comparing results on Mistral-7B (Regen-M) and Llama-2-7B (Regen-L) generations (TPR@5%FPR).
(a) Blahut–Arimoto weights without sampling on Llama-2-7b logits.
(b) Blahut–Arimoto weights with sampling on Llama-2-7b logits.
@@
(a) Blahut–Arimoto weights on Pubmed generated with Llama-2-7b.
(b) Blahut–Arimoto weights on Pubmed generated with Llama-2-7b with Phi in the ensemble.
(c) Blahut–Arimoto weights on Pubmed generated with Llama-2-7b with our Unigram model in the ensemble
Figure 2: Comparison of Blahut–Arimoto weights on the Pubmed Regen-L dataset for different RSA  configurations.
Appendix D Ghostbuster Results
When it comes to the Ghostbuster dataset, all methods perform extremely well, with RSA-4 lagging a bit behind. However, unlike for Binoculars, adding Phi into the mix is of great help. We believe this is because ChatGPT’s distribution is closer to Phi’s than the other models in our mix. Adding one more model to our RSA-4 thus enables us to approach the generator’s distribution better.
Method Writing Prompts Reuter Essay Average
PPL TowerBase-7b 0.973 0.961 0.990 0.975
PPL TowerBase-13b 0.976 0.964 0.991 0.977
PPL Llama-2-7b-chat 0.989 0.994 0.994 0.992
PPL Llama-2-7b 0.977 0.971 0.991 0.980
PPL Phi-3-mini-4k-instruct 0.993 0.990 0.995 0.993
$q^{\star}$ log-probs 0.985 0.984 0.994 0.988
$q^{\star}$ 5 (+phi) log-probs 0.992 0.990 0.995 0.992
Binoculars Score 0.990 0.993 0.996 0.993
RSA(avg) 0.946 0.971 0.911 0.943
RSA(max) 0.975 0.979 0.974 0.976
RSA-5 (+phi, avg) 0.975 0.986 0.966 0.976
RSA-5 (+phi, max) 0.993 0.996 0.995 0.995
Table 7: AUC Scores on the Ghostbuster dataset.
Method Writing Prompts Reuter Essay Average
PPL TowerBase-7b 0.910 0.789 0.961 0.887
PPL TowerBase-13b 0.924 0.810 0.971 0.902
PPL Llama-2-7b-chat 0.967 0.983 0.977 0.976
PPL Llama-2-7b 0.924 0.854 0.968 0.915
PPL Phi-3-mini-4k-instruct 0.974 0.962 0.987 0.974
$q^{\star}$ log-probs 0.953 0.922 0.982 0.952
$q^{\star}$ 5 (+phi) log-probs 0.974 0.955 0.987 0.972
Binoculars Score 0.960 0.963 0.987 0.970
RSA(avg) 0.737 0.843 0.558 0.713
RSA(max) 0.887 0.886 0.882 0.885
RSA-5 (+phi, avg) 0.899 0.933 0.832 0.888
RSA-5 (+phi, max) 0.979 0.983 0.986 0.983
Table 8: TPR@5%FPR on the Ghostbuster dataset.
Appendix E arxiv-cs.cl
As mentioned in Section 4.3, both the baseline Binoculars and our RSA method perform poorly on this dataset and are unable to effectively differentiate between generated and original texts. As illustrated in Figure 3, the scores computed by both methods on human and artificial texts largely overlap, resulting in a detection performance that is close to a random decision.
Figure 3: Histograms of scores on the scholarly dataset.
Appendix F M4
Table 9 displays all our results on the M4 multilingual datasets we used. We notice that RSAperforms better on Chinese Davinci than Chinese ChatGPT but it is the opposite when it comes to Russian and Bulgarian. On those two languages, performances quickly drop when looking at the TPR@5%FPR.
On Figure 4, we compare the Arimoto weights on texts from the same domain (news articles) but different languages. The English one favours Llama-2-7b very heavily, while the bulgarian version gives more credit to the TowerBase models we use, which is what we expect since those models have seen more multilingual data during training.
AUC Arabic German Chinese Russian Bulgarian
Chatgpt ChatGPT ChatGPT DaVinci ChatGPT DaVinci ChatGPT DaVinci
PPL (oracle) 0.918 0.693 0.983 0.946 0.695 0.391 0.932 0.449
BinocularsScore 0.686 0.914 0.982 0.979 0.674 0.554 0.742 0.659
RSA(avg) 0.909 0.890 0.930 0.980 0.737 0.324 0.974 0.556
RSA(max) 0.912 0.885 0.909 0.964 0.688 0.259 0.956 0.642
TPR @ 5% FPR
PPL (oracle) 0.587 0.052 0.945 0.862 0.199 0.011 0.719 0.016
BinocularsScore 0.129 0.679 0.929 0.957 0.085 0.038 0.202 0.117
RSA(avg) 0.564 0.603 0.609 0.951 0.220 0.016 0.898 0.064
RSA(max) 0.558 0.571 0.498 0.864 0.139 0.013 0.789 0.099
Table 9: Detailed results for the M4 dataset.
(a) Blahut–Arimoto weights when looking at the llama-2-7b generated cc-news dataset.
(b) Blahut–Arimoto weights when looking at Bulgarian ChatGPT generations.
Figure 4: Comparison of Blahut–Arimoto weights between English and Bulgarian.

### v2

@@ 6 Summary and Discussion
Our MOSAIC method effectively harnesses the ensemble’s strength, achieving great results across datasets and languages, eliminating the need to find the best detector while offering a scalable solution that can incorporate future models. However, it is currently computationally costly, as each model must run on the text (while not the goal of this work, some improvements are proposed in Section D). We also need to develop information-theoretic tools to select the most useful models and filter out less effective ones, as shown with the unigram model. Similarly, the proximity of models needs further study, as we suspect our detectors are too far from the fine-tuned model’s generating distribution used to create the academic dataset, hence why it evades detection.
@@ 7 Ethics Statement
8 Reproducibility Statement
All our experiments use public benchmarks and open-source code. The additional corpora created for this study, as well as all the code used to implement the MOSAIC algorithm, will also be openly released.
Acknowledgements
This work was performed using HPC resources from GENCI–IDRIS (Grant 2023-AD011014903).
@@ Appendix A Proof of Proposition 1
where the minimum is taken over all the possible distributions $q\in\mathcal{P}(\Omega)$, representing the expected value of regret of $q$ w.r.t. the worst-case distribution over $\mu\in\mathcal{P}(\mathcal{M})$. Notice that this is equivalent to the average worst-case regret Barron et al. [1998], Silva and Piantanida [2022]. The equality in (12) holds by noticing that
@@
is convex for every $(\mu,q)$, respectively. Then, by classical min-max theorem von Neumann [1928], we have that (15) holds.
Finally, it remains to show that:
@@
We begin by showing that:
@@ B.1 Algorithm
Now if we denote the cardinality $|\mathbb{M}|=M$, $|\Omega|=N$, then $p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$ is an $M\times N$ matrix, which we denote the $i$-th row, $j$-th column entry by $w_{ij}$. For the case of channel capacity, the algorithm was introduced in Arimoto [1972], Blahut [1972] to solve (20). They both found the following expression for the capacity of a discrete channel with channel law $w_{ij}$:
@@
Then, using the theory of optimization, specifically coordinate descent, it has been shown that the sequence indeed converges to the required maximum. That is,
@@ B.2 Computational complexity
Appendix C Binocular Scores
@@
$B(\mathbf{w})\triangleq\frac{\displaystyle\displaystyle\sum_{t=1}^{T}\sum_{y_{t}\in\Omega}\mathbbm{1}[y_{t}=w_{t}]\mathcal{L}_{\theta}(y_{t}|m,\mathbf{w}_{<t})}{\displaystyle\sum_{t=1}^{T}\sum_{y_{t}\in\Omega}\displaystyle p_{\theta}(y_{t}|m^{\prime},\mathbf{w}_{<t})\mathcal{L}_{\theta}(y_{t}|m,\mathbf{w}_{<t})},$ (23)
@@
where the choices of LLMs $p_{\theta}(\cdot|m,\mathbf{w}_{<t})\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ and $p_{\theta}(\cdot|m^{\prime},\mathbf{w}_{<t})$ $\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ are critical for performance and have to be optimized empirically. Indeed, this represents the main weakness of this score, since in practice the best choice for the best pair of LLMs $(m,m^{\prime})$ may not be distribution-free.
Appendix D Complexity Improvements
Our algorithm currently processes each text in approximately 10 seconds on NVIDIA 32G V100 GPUs. Runtime optimization is an area that should be improved in future work. Below, we outline limitations of our system and propose potential improvements : In MOSAIC, the texts are processed one-by-one by the LLMs. Each model is loaded onto a separate GPU, and the logits are moved to a central device for performing operations such as Blahut-Arimoto, perplexity, and cross-entropy calculations, after which the final score is computed. This setup has several inefficiencies. For instance, transferring logits to a central device introduces a significant bottleneck. Additionally, while calculations are performed on one GPU, the remaining ones remain idle, resulting in suboptimal use of resources.
A more efficient method would involve computing the logits for all texts in parallel, storing them across different GPUs, and performing subsequent calculations concurrently. An even more streamlined solution would involve loading all models onto a single GPU using quantized or distilled versions, thus eliminating the need to transfer logits across devices.
While these optimizations are promising, they have not been implemented in this work, as we focus on the algorithmic methodology rather than runtime efficiency.
Appendix E Robustness results
Table 5 is a more complete version of Table 3, where we also include the details of all individual detectors based on just one model.
Table 5: Artificial text detection performance of detectors built with a fixed set of 4 models. Detection may involve running 1, 2, or 4 models. AUROC scores.
RAID Binoculars Ghostbuster M4 (multilingual) Scho. Avg
Pubmed CNN CCnews Reddit Reuter Essay Ara Bul Ger Rus
1 model
PPL based detectors with …
Tower-7b 0.709 0.798 0.557 0.582 0.973 0.961 0.990 0.882 0.839 0.639 0.523 0.503 0.746
Tower-13b 0.705 0.823 0.561 0.578 0.976 0.964 0.991 0.879 0.837 0.598 0.514 0.522 0.746
Llama-2-7b 0.723 0.786 0.544 0.556 0.977 0.971 0.991 0.888 0.896 0.597 0.626 0.520 0.756
Llama-2-7b-chat 0.769 0.817 0.598 0.617 0.989 0.994 0.994 0.918 0.932 0.693 0.695 0.520 0.795
DetectGPT-based detectors with …
Tower-7b 0.481 0.448 0.551 0.473 0.635 0.578 0.888 0.579 0.667 0.299 0.630 0.476 0.559
Tower-13b 0.449 0.489 0.555 0.489 0.616 0.555 0.892 0.578 0.642 0.267 0.623 0.480 0.553
Llama-2-7b 0.498 0.430 0.530 0.451 0.674 0.650 0.925 0.606 0.690 0.278 0.656 0.477 0.572
Llama-2-7b-chat 0.598 0.450 0.559 0.476 0.852 0.810 0.949 0.735 0.758 0.391 0.684 0.485 0.646
FastDetectGPT-based detectors with …
Tower-7b 0.825 0.997 0.997 0.965 0.914 0.834 0.849 0.973 0.979 0.896 0.781 0.531 0.878
Tower-13b 0.834 0.999 0.995 0.975 0.878 0.886 0.818 0.985 0.988 0.832 0.816 0.517 0.877
Llama-2-7b 0.810 0.993 0.994 0.959 0.647 0.616 0.463 0.954 0.990 0.893 0.879 0.511 0.809
Llama-2-7b-chat 0.744 0.955 0.877 0.896 0.324 0.650 0.127 0.870 0.772 0.644 0.591 0.395 0.654
2 models
Bino-best 0.803 0.989 0.994 0.973 0.677 0.663 0.481 0.897 0.959 0.860 0.492 0.381 0.764
4 models
avg PPL 0.730 0.808 0.566 0.584 0.980 0.980 0.992 0.897 0.887 0.638 0.598 0.516 0.765
$q^{\star}$ (log-probs) 0.746 0.807 0.566 0.579 0.985 0.984 0.994 0.893 0.901 0.640 0.592 0.509 0.766
MOSAIC-4 (avg) 0.850 0.992 0.993 0.971 0.946 0.971 0.911 0.909 0.974 0.890 0.737 0.421 0.880
MOSAIC-4 (unif) 0.844 0.992 0.995 0.975 0.920 0.951 0.876 0.909 0.974 0.893 0.745 0.416 0.874
Table 6 displays TPR @ 5% FPR obtained when running our system on the RAID test dataset. It can be seen that adding repetition penalty (the w/ r_p columns) makes our results drop significantly. Note that, as no labels are provided, this is a different metric from the other tables.
Table 6: MOSAIC performance under different generator configurations on the RAID test dataset. Scores are TPR@5%FPR.
Greedy Greedy w/o r_p Greedy w/ r_p Sampling Sampling w/o r_p Sampling w/ r_p Repetition Penalty No Repetition Penalty All
MOSAIC-4 0.902 0.952 0.810 0.603 0.785 0.269 0.540 0.868 0.752
MOSAIC-5 0.884 0.927 0.806 0.606 0.799 0.252 0.529 0.863 0.745
Table 7: MOSAIC under different adversarial attacks. Scores are TPR@5%FPR.
All White Space Upper Lower Synonym Miss- spelling Para- phrase Number Shuffling Add Paragraphs Homo- glyph Article Deletion Change Spelling Zero Width Space
MOSAIC-4 0.693 0.675 0.686 0.285 0.725 0.719 0.713 0.745 0.866 0.708 0.729 0.714
MOSAIC-5 0.694 0.670 0.665 0.227 0.717 0.703 0.697 0.733 0.902 0.695 0.722 0.855
Figure 2 represents how Arimoto weights evolve when looking at another language. As the TowerBasemodels have been trained on more multilingual data, they have more importance when looking at Bulgarian text.
(a) Blahut–Arimoto weights for Llama-2-7b generations in English (CC_news dataset).
(b) Blahut–Arimoto weights for ChatGPT generations in Bulgarian.
Figure 2: Comparison of Blahut–Arimoto weights between English (CC_news) and Bulgarian (M4).
Figure 3: Scores obtained on the Academic dataset, MOSAIC on the left and Binoculars on the right, for both scoring methods, generated and original texts are indistinguishable
Appendix F Including the generator in the ensemble
In this section, we report in Table 8 the TPR @5% FPR corresponding to the AUROC scores in Table 4. We also display in Figure 4 the changes in Blahut-Arimoto weights when simulating the effect of nucleus-sampling on the logits computed by Llama-2-7b.
Table 8: Identification results with varying generators for the Binoculars dataset: the original version (Llama-13b and Falcon), and our regenerated corpus with Llama-2-7B (Llama) and Mistral-7B (Mist.) (TPR@5%FPR).
Pubmed CNN CC_news Avg.
Orig. Llama Mist. Orig. Llama Mist. Orig. Llama Mist.
@@ Appendix F Including the generator in the ensemble
MOSAIC-4 (avg) 0.963 0.615 0.814 0.971 1.000 0.898 0.868 0.552 0.730 0.823
(a) Blahut–Arimoto weights for Pubmed regenerated with Llama-2-7b.
(b) Blahut–Arimoto weights for Pubmed regenerated with Llama-2-7b. In the detector, the logits for Llama-2-7b are modified to simulate the effect of nucleus sampling.
@@
Appendix G Increasing the size of the ensemble
In this section, we add a Phi-3 model and look at the difference in results, tables 9 and Arimoto weights, Figure 5, as well as when adding a Unigram model on Figure 6.
(a) Weights for the essay dataset when phi is in the ensemble.
(b) Weights for the essay dataset for the ensemble without phi.
Figure 5: Comparison of Blahut–Arimoto weights between English (CC_news) and Bulgarian (M4).
Table 9: AUROC Scores on Ghostbuster and Binoculars datasets.
Method Reddit Reuter Essay Pubmed CNN CC_news Avg.
PPL TowerBase-7b 0.973 0.961 0.990 0.709 0.798 0.557 0.831
PPL TowerBase-13b 0.976 0.964 0.991 0.705 0.823 0.561 0.837
PPL Llama-2-7b-chat 0.989 0.994 0.994 0.769 0.817 0.598 0.860
PPL Llama-2-7b 0.977 0.971 0.991 0.723 0.786 0.544 0.832
PPL Phi-3-mini-4k-instruct 0.993 0.990 0.995 0.586 0.997 0.495 0.843
$q^{\star}$ log-probs 0.985 0.984 0.994 0.807 0.566 0.579 0.819
$q^{\star}$ (+phi) log-probs 0.992 0.990 0.995 0.632 1.000 0.561 0.862
MOSAIC-4 (avg) 0.946 0.971 0.911 0.992 0.993 0.971 0.964
MOSAIC-5 (+phi, avg) 0.975 0.986 0.966 0.992 0.988 0.950 0.976
MOSAIC-5 (+1gram, avg) - - - 0.986 0.960 0.898
(a) Weights for the pubmed dataset without the unigram model.
(b) Weights for the pubmed dataset when the unigram is in the ensemble.
Figure 6: Comparison of Blahut–Arimoto weights when adding the unigram to our ensemble.